Show that the differential equation

2xz + q^2 = x(xp + yq)

has a complete integral


z+ a^2x = axy + bx^2

and deduce that

x(y + hx)^2 = 4(z - kx^2)

is also a complete integral.
1
Expert's answer
2020-11-24T16:54:53-0500

Let "F(x,y,z,p,q)=2xz+q^2-x(xp+yq)=0."Then Charpit’s equations are


"\dfrac{dx}{-x^2 }=\dfrac{dy}{2q-xy}=\dfrac{dz}{-px^2+2q^2-qxy}="

"=\dfrac{dp}{-2z+yq}=\dfrac{dq}{-qx}"

From the first and last equations, we get


"\dfrac{dx}{x }=\dfrac{dq}{q }"

Integrate


"\ln|x|+\ln a=\ln|q|"

"q=ax, a=const."

Then the equation "F(x, y, z, p, q)=0" gives


"2xz+a^2x^2-x(xp+ayx)=0"

"p=\dfrac{2z+a^2x-ayx}{x}"

Hence the relation "dz=pdx+qdy" leads to


"dz=\dfrac{2z+a^2x-ayx}{x}dx+axdy=>"

"=>d(\dfrac{z}{x^2})+d(\dfrac{a^2}{x})=ad(\dfrac{y}{x})"

which, on integration, gives


"\dfrac{z}{x^2}+\dfrac{a^2}{x}=a\dfrac{y}{x}+b"

"z+a^2x=axy+bx^2"

which is a complete integral, "b" being a constant.


Now we show that

"x(y + cx)^2 = 4(z − dx^2)\ \ \ \ \ \ \ \ \ (1)"

is also a complete integral.

Let us consider the curve Γ : "y=0, z=\dfrac{c^2x^3+4dx^2}{4}" on the surface "(1)."

At the intersections of "z+a^2x=axy+bx^2" and the curve Γ we have


"c^2x^2 + 4(d \u2212 b)x + 4a^2 = 0"

which has equal roots when "b=d\pm ac." Taking "b=d+ac," the subsystem has the equation "z+a^2x=axy+(d+ac)x^2," i.e.

"a^2x-x(cx+y)a+(z-dx^2)=0"

which has the envelope


"x^2(cx+y)^2=4x(z-dx^2),"

i.e.


"x(cx+y)^2=4(z-dx^2)"

This is, therefore, a complete integral.